For Company the following information is given

Beginning Inventory 115,000 units Cost = $2,156,250

Produced during period 140,000 units Cost = $2,915,700

They sold 201,000 units during the year, using the average cost inventory method what is the COGS?

A. $3,768,750

B. $3,997,890

C. $4,186,112

D. $6,577,174
Answer #1

Total cost = 2,156,250 + 2,915,700 = 5,071,950

Totak units available for sale = 115,000 + 140,000 = 255,000

Average cost per unit = 5,071,950/255,000

= 19.89 per unit

Cost of Goods Sold = 201,000 * 19.89

= 3,997,890

Option B is the answer